OpenStudy (anonymous):

Find the general solution of: sin^2 x= 3 cos^2 x

sam (.sam.):

Where you got stucked?

OpenStudy (anonymous):

I have the general solution of \[x=\frac{ \pi }{ 3 } \pm 2\pi n\]

OpenStudy (anonymous):

Just wanted to make sure I was right

sam (.sam.):

Should be \(x=\frac{ \pi }{ 3 } \pm \pi n\)

OpenStudy (anonymous):

Ok, thank you